Find the energy equivalent of one atomic mass unit, first in Joules and then in MeV. Using this, express the mass defect of 16 (8) O in MeV/c^2.

Grade:11

1 Answers

vasu
7 years ago
I amu=1.66×1027 kg c = speed of light = 3×108 m/sE = mc2E= 1.66×1027×( 3×108)2E = 1.49×10−10 joule1 MeV= 1.6×10−13 JE = 1.49×10−101.6×10−13= 931.6 MeVO168mass no = 16energy for 16 amu = 16×931.6 = 14905.6 MeVmass defect = energyc2 = 14905.6 MeV/c2